OpenStudy (anonymous):

Emilio's age is four years more than half the age of his sister. Their sum is 28.how old is Emilio? 12 16 20 or 5

OpenStudy (ankit042):

let us assume Emillo age as x and the sister age as y now x-4 = y/2 and x+y =28

OpenStudy (ankit042):

Now you can solve the equations to find the value of x and y

OpenStudy (anonymous):

Thanks

OpenStudy (ankit042):

np :)